OpenStudy (anonymous):

Express the complex number in trigonometric form. -6i

OpenStudy (anonymous):

\(-i\) means \(\theta =3\pi/2\) and \(6\) means \(r=6\).

OpenStudy (anonymous):

@wio so how would i put that into one statement

OpenStudy (anonymous):

6(cos 0° + i sin 0°) 6(cos 270° + i sin 270°) 6(cos 180° + i sin 180°) 6(cos 90° + i sin 90°) those are the answer choices and their formats

OpenStudy (anonymous):

Are you phoning this one in? What class is this?

OpenStudy (anonymous):

it's pre-calc I read through the lesson, but i'm still confused as to how to get from point a (the number) to point b (the final form)

OpenStudy (anonymous):

Basically, \((a,b)\) are the Cartesian coordinates, and \((r,\theta)\) are the polar coordinates. \[ a+bi = re^{i\theta} = r[\cos\theta+i\sin\theta] \]

OpenStudy (anonymous):

In our case, we have \((a,b) = (0,-6)\). We want \((r,\theta)\).

OpenStudy (anonymous):

\[ r= \sqrt{a^2+b^2} \\ \theta = \arctan\left(\frac ba\right) \]

OpenStudy (anonymous):

In this case, since \(a=0\), we have to remember that \((0,-6)\) it directly under the origin, that's a trivial angle to find.

OpenStudy (anonymous):



OpenStudy (anonymous):

alternatively, use the fact that \(\cos(\theta) = 0\) and \(i\sin(\theta) = -1\).

OpenStudy (anonymous):

How would you find theta if a=0 then, is there a special case for it?

OpenStudy (anonymous):

When \(a=0\), that means \(\cos\theta = 0\). So \(\theta\) must be \(\pi/2\) or \(3\pi/2\).

OpenStudy (anonymous):

and how can you tell which one it is?

OpenStudy (anonymous):

\(b>0\implies \theta = \pi/2\). \(b<0 \implies \theta =3\pi/2\). Did you read what I wrote and drew before?

OpenStudy (anonymous):

yes, so it would be 3pi/2

OpenStudy (anonymous):

which i guess makes the answer B? 6[cos270+isin270]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thank you for your help!